A ball is thrown vertically upwards from ground it crosses a point 25 m  twice at an interview of 4 s find velocity at which ball was thrown 

Grade:11

1 Answers

Vikas TU
7 years ago
AT top
0  =  u – at
0 = u -10*2
u = 20 m/s (t = 2s)
u becomes final velocity for the ground projection thus,
400 = u^2 – 2*-10*25
u = 30 m/s
This would be gthe velocity.